OpenStudy (anonymous):

Please Help will give medal! Frankie can build a speaker box in 120 minutes. Tisha can build a speaker box 40 minutes faster. If they work together, how many minutes does it take them to build the speaker box?

OpenStudy (anonymous):

I think it's B

OpenStudy (♪chibiterasu):

There are no options but ok: x/120 +x/40 =1 x +3x =120 4x=120 x=30 minutes \[\frac{ x }{ 120 } + \frac{ x }{ 40 } = 1\] x + 3x = 120 4x = 120 x = ?

OpenStudy (anonymous):

x = 30

OpenStudy (♪chibiterasu):

Yeah.

OpenStudy (anonymous):

It's not one of the choices, I'm confused.

OpenStudy (anonymous):

anyone? please help?

OpenStudy (anonymous):

@Hi-xx @hughfuve @pandaecat @bruno102 @gabbimanges17 @The_Yee_Dinosaur @ctwatson @ximevazq @dwoodruff

hero (hero):

1. The amount of time it will take when they work together will be less than the time it takes Tisha to do it alone. 2. To find the time it takes for Frankie and Tisha to build the speaker together, let F = the time it takes Frankie to build the speaker alone T = the time it takes Tisha to build the speaker alone Then input those values into the following expression: \[\dfrac{F \times T}{F + T}\] and then evaluate the expression.

OpenStudy (anonymous):

It's 30 but that's not the answer the choices are 40 minutes 48 minutes 80 minutes 240 minutes

hero (hero):

If it were me, I'd pick the lowest amount. 40 could be a typo.

OpenStudy (anonymous):

Alright I'll try...

OpenStudy (anonymous):

It was wrong. :(

hero (hero):

If that's wrong, then the other choices are more wrong.

hero (hero):

30 is definitely the right amount of minutes.